A body constrained to move along the z-axis of a coordinate system is subject to a constant force F given by F = −i + 2 j + 3 k N where i, j, k are unit vectors along the x-, y- and z-axis of the system respectively. What is the work done by this force in moving the body a distance of 4 m along the z-axis?
1
Expert's answer
2019-10-22T10:20:10-0400

The work done


"W=\int F_zdz=\int 3dz=3z=3\:N\times 4\:m=12\: J"